# Generalists vs. Specialists: Evaluating Large Language Models for Urdu

### v2

@@ Abstract
In this paper, we compare general-purpose pretrained models, GPT-4-Turbo and Llama-3-8b-Instruct with special-purpose models fine-tuned on specific tasks, XLM-Roberta-large, mT5-large, and Llama-3-8b-Instruct. We focus on seven classification and six generation tasks to evaluate the performance of these models on Urdu language. Urdu has 70 million native speakers, yet it remains underrepresented in Natural Language Processing (NLP). Despite the frequent advancements in Large Language Models (LLMs), their performance in low-resource languages, including Urdu, still needs to be explored. We also conduct a human evaluation for the generation tasks and compare the results with the evaluations performed by GPT-4-Turbo and Llama-3-8b-Instruct. We find that special-purpose models consistently outperform general-purpose models across various tasks. We also find that the evaluation done by GPT-4-Turbo for generation tasks aligns more closely with human evaluation compared to the evaluation by Llama-3-8b-Instruct. This paper contributes to the NLP community by providing insights into the effectiveness of general and specific-purpose LLMs for low-resource languages.
@@
Samee Arif1, Abdul Hameed Azeemi1, Agha Ali Raza1, Awais Athar2, {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk, awais@ebi.ac.uk 1Lahore University of Management Sciences, 2EMBL European Bioinformatics Institutes
@@ 1 Introduction
    Generalists: General-purpose models capable of performing a wide variety of tasks. We will use GPT-4-Turbo (abbreviated as GPT-4) trained on dataset up to Dec 2023 and Llama-3-8b-Instruct (abbreviated as Llama) as the generalist models.
@@
    Specialists: Special-purpose models fine-tuned to perform specific tasks. We use XLM-Roberta-large (abbreviated as XLM-R) (Conneau et al., 2020), mT5-large (abbreviated as mT5) (Xue et al., 2021), and a fine-tuned version of Llama-3-8b-Instruct (abbreviated as Llama-FT) as the specialist models.
@@
PoS Tagging Translation (ur-en)
NER Tagging
@@
    We conduct a human evaluation of the outputs from the generation tasks and compare these results with the automated evaluations performed by GPT-4 and Llama.
The code, model outputs, and the human and LLM evaluations are publicly available on GitHub [^1] https://github.com/ulrs0/Generalists-vs-Specialists.
@@ 2 Related Work
In recent years there has been a growing interest in the performance of LLMs across various languages. The MEGA benchmark (Ahuja et al., 2023) evaluates 16 NLP datasets across 70 languages. They compare the performance of BLOOMZ, GPT models, and State of the Art (SOTA) non-autoregressive models. MEGAVERSE Ahuja et al. (2024) builds on top of the MEGA benchmark and evaluates the non-English capabilities of GPT-3.5-Turbo, GPT-4, PaLM2, Gemini-Pro, Mistral, Llama2, and Gemma. IndicGenBench (Singh et al., 2024) evaluates LLMs on user-facing generation tasks across a set of 29 Indic languages. They perform evaluation on both proprietary and open-source LLMs including GP-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM, and Llama. They cover the following tasks: cross-lingual summarization, machine translation, and cross-lingual question-answering. Zhao et al., 2024 conduct an evaluation of Llama’s response quality based on LLM-Eval (Zhang et al., 2023), a benchmark comprising instruction tasks from 17 categories. Mujadia et al., 2024 presents a comprehensive translation evaluation using Llama for English and Indian languages. They found that LLM-based evaluator achieves a comparable score with human judgement.
Khondaker et al., 2023a presents a in-depth evaluaton of BLOOMZ, ChatGPT and specialist models AraT5 and MARBERTv2. They evaluated these models for Arabic on 44 distinct language understanding and generation tasks. They also present a comparison between the human evaluation and GPT-4 evaluation. Additionally, they observed that the English prompt works better than the Arabic prompt. Abdelali et al., 2024 provides a benchmark of LLMs against SOTA models for Arabic NLP. They evaluated GPT-3.5-Turbo, GPT-4, BLOOMZ, and Jais-13b-Chat across 33 tasks using 61 publicly available datasets, resulting in 330+ experimental setups. They observed that SOTA models generally outperform LLMs in zero-shot learning, larger models with few-shot learning techniques significantly reduce the performance gaps.
Existing multilingual NLP benchmarks often lack extensive language-specific evaluation and comparison against task-specific models. Tahir et al., 2024 address this gap by evaluating GPT-3.5-Turbo, Llama2-7B-Chat, and Bloomz (3B and 7B1) across 14 tasks using 15 Urdu datasets in a zero-shot setting, comparing their performance against task-specific models. Their findings reveal that task-specific models (Support Vector Machine (SVM), Decision Tree, and m-BERT, etc) generally outperform the mentioned LLMs in Urdu NLP tasks with zero-shot learning. However, they do not perform few-shot prompting, chain-of-thought prompting, or LLM fine-tuning.
@@ 3.1 Classification
Abuse Detection. We use the dataset by Akhter et al., 2020 which has 2,171 entries and the dataset by Amjad et al., 2022 which consists of 3,502 entries.
@@
Fake News Detection. We use the fake news dataset by Amjad et al., 2020 which is a mixture of real and translated data. It comprises 1,300 labeled news articles.
@@
Part-of-Speech Tagging. For this task, we use the Universal Dependencies (Nivre et al., 2020) dataset, which consists of 5,130 sentences, annotated with a part-of-speech tag for every word. For GPT-4 and Llama, we wrap the word for which we want to predict the part-of-speech tag in <hl> tags. The structure of the data is given in Figure 1.
@@
Named Entity Recognition. We use the dataset [^4] https://huggingface.co/datasets/mirfan899/urdu-ner available on Hugging Face which consists of 33,748 sentences, annotated with a NER tag for every word. For GPT-4 and Llama, we applied the same data structuring method as done in the part-of-speech tagging task.
@@ 3.2 Generation
Given our resource constraints, we could not set the maximum sequence length for model training to the longest entries in each dataset. Instead, we filter the datasets to include only those entries that fell within a manageable maximum length. This approach allowed us to optimize the training process and ensure efficient use of computational resources while still maintaining a representative sample of the data.
Question-Answering. We use three datasets for question-answering: (1) UQA (Arif et al., 2024), consisting of 88,829 answerable questions; (2) UQuAD [^5] https://github.com/ahsanfarooqui/UQuAD---Urdu-Question-Answer-Dataset/tree/main, containing 139 questions; and (3) Wiki-UQA [^6] https://huggingface.co/datasets/uqa/Wiki-UQA, a manually generated dataset from Wikipedia articles, comprising 210 questions.
@@ 4.1 Experimental Design
We utilize a controlled test set consisting of min(1000, len(dataset["test"])) samples for each task, ensuring that the evaluation is both comprehensive and cost-efficient. We ensure that the test dataset is balanced by having an equal representation of different classes within each task. We use Macro-$F_{1}$ Score to evaluate all the classification tasks, SQuAD (Rajpurkar et al., 2018) Rajpurkar et al. (2016) $F_{1}$ to evaluate question-answering, SacreBLEU (Post, 2018) for paraphrasing, transliteration and translation and ROUGE-L (with word-level tokenization) Lin (2004) for summarization.
@@
To evaluate the performance of the generalist models, we design a series of experiments. We use GPT-4-Turbo and Llama-3-8b-Instruct as our generalist models. GPT-4-Turbo is chosen due its top performance on the LMSYS chatbot arena (Chiang et al., 2024) on of March 1st, 2024 when we started our research. Our experimental setup is as follows: For GPT-4 We conduct experiments under zero-shot, three-shot, and six-shot settings for both generation and classification tasks. The examples for the three-shot and six-shot scenarios are selected from the training dataset of each task. Specifically, the examples are carefully selected by a human expert to ensure a representative and balanced sample, avoiding any unintended bias in the selection process. We also conduct experiments with chain of thought (CoT) reasoning for classification tasks in the six-shot setting. We create CoT reasoning for the six selected examples from the training dataset. These examples along with their CoT is given as a few-shot prompt to the model. Figure 2 presents an example of generated CoT reasoning. For all the evaluations, the temperature and nucleus sampling for GPT-4 is set to 1.0, which is the default for the GPT API.
@@
For Llama, we only perform zero-shot evaluations since it is an instruct model (not fine-tuned for chat). This helps us establish a baseline performance for the generalist models and aids in comparison with the fine-tuned specialist Llama-FT. For Llama, the temperature is set to 0.6 and nucleus sampling is set to 0.9, i.e., the default values in the original code-base for Llama [^8] https://github.com/meta-llama/llama3/blob/main/example_chat_completion.py.
@@ 4.2 Prompt Design
Khondaker et al., 2023b observe in their ChatGPT evaluation for Arabic that an English prompt performs better than the Arabic one. Therefore, following this study, we decide to use English prompts for our evaluations of Urdu tasks as well. In the prompt templates, the following placeholders are used:
@@ 5.1 Classification
GPT-4-Turbo Llama-3-8b-Instruct XLM-R-large Llama-3-8b-Instruct-FT
Task 0 3 6 CoT
Sentiment Analysis 90.98 91.17 94.90 94.60 87.88 92.90 95.30
Abuse Detection 86.27 89.01 88.71 87.62 44.73 90.92 89.15
Sarcasm Detection 58.17 69.18 66.56 65.47 44.94 84.37 81.48
Fake News Detection 78.88 75.95 78.14 76.45 66.36 84.99 72.14
Topic Classification 76.05 73.54 74.57 70.43 54.08 84.49 84.74
PoS Tagging 53.31 51.51 54.17 54.61 25.80 65.41 67.55
NER Tagging 61.96 62.18 62.98 63.95 42.13 70.41 90.90
Table 2: We report Macro-$F_{1}$ score for each classification task. GPT-4-Turbo is evaluated in 0-shot, 3-shot, 6-shot and 6-shot with chain-of-thought settings. In Llama-3-8b-FT, FT stands for fine-tuned.
We present the evaluation of the generalists (GPT-4, Llama) and the specialists (Llama-FT, XLM-R) for the classification tasks in Table 2. We observe that Llama-FT (fine-tuned) model achieves the highest scores in four tasks (sentiment analysis, topic classification, PoS tagging, and NER tagging), while XLM-R outperforms other models in three tasks (abuse detection, sarcasm detection, and fake news detection). GPT-4 does not perform better than Llama-FT and XLM-R. For certain tasks like NER and PoS tagging, chain-of-thought (CoT) reasoning leads to a better Macro-$F_{1}$ score. We now discuss the performance on each classification task in detail.
Sentiment Analysis. For sentiment analysis, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 95.30, outperforming XLM-R and GPT-4-Turbo. GPT-4’s performance improves with more shots, reaching a Macro-$F_{1}$ of 94.90 with 6-shot learning. The CoT setting provides a slight increase to 94.60. The small difference in the Macro-$F_{1}$ between Llama-FT and GPT-4 indicates the effectiveness of GPT-4 for Urdu sentiment analysis without requiring task-specific fine-tuning.
@@
NER Tagging. For NER tagging, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 90.90, significantly outperforming XLM-R and GPT-4. The best score for GPT-4 is 63.95 with CoT reasoning. We uncover the differences in accuracy for various entities for GPT-4 and Llama-FT in Figure 8. We observe a drop in the recognition accuracy of Person entities by GPT-4, with 95 correctly recognized compared to 139 by Llama-FT. We notice a similar trend for Organization entities, with 70 classified correctly compared to 134 by Llama-FT. This suggests that NER fine-tuning on a specialized Urdu dataset improves the model’s ability to recognize Person and Organization entities in Urdu text.
Abuse Detection. In the task of abuse detection, XLM-R leads with a Macro-$F_{1}$ of 90.92, followed closely by Llama-3-8b-FT at 89.15. GPT-4’s performance peaks at 89.01 with 3-shot learning and slightly decreases with CoT reasoning.
Sarcasm Detection. XLM-R performs best in sarcasm detection with a Macro-$F_{1}$ score of 84.37. Llama-FT also shows strong performance with 81.48. GPT-4’s best performance is observed at 69.18 with 3-shot learning. The large difference between the Macro-$F_{1}$ scores of generalist and specialist models indicates that sarcasm detection in Urdu can be challenging without task-specific fine-tuning.
Fake News Detection. For fake news detection, XLM-R achieves the highest score with a Macro-$F_{1}$ of 84.99, while Llama-FT follows with 72.14. GPT-4’s performance is relatively consistent across different shot settings, with its highest score being 78.88 in the 0-shot setting. The large margin by XLM-R indicates its effectiveness in discerning fake news.
@@
PoS Tagging. In PoS tagging, Llama-FT leads with a Macro-$F_{1}$ of 67.55, followed by XLM-R at 65.41. GPT-4’s highest score is 53.31 in the 0-shot setting. To understand the significant difference in Macro-$F_{1}$ between GPT-4 and Llama-FT, we study the individual class performance (Figure 9). We find that GPT-4 struggles with correctly tagging proper nouns, pronouns, and auxiliaries, while Llama-FT is able to identify most of them correctly, suggesting that task-specific fine-tuning improves tagging performance for these parts of speech.
Metric GPT-4-Turbo Llama-3-8b-Instruct mT5-large Llama-3-8b-Instruct-FT
Task 0 3 6
Question-Answering SQuAD-$F_{1}$ 66.28 72.55 72.78 69.89 69.66 70.42
Summarization ROUGE-L 22.54 23.35 23.34 25.76 30.72 31.01
Paraphrasing SacreBLEU 3.59 3.92 4.01 2.42 11.98 10.17
Transliteration SacreBLEU 30.93 32.38 32.09 12.37 40.23 37.95
Translation (en-ur) SacreBLEU 12.07 12.62 11.59 5.14 18.35 14.63
Translation (ur-en) SacreBLEU 16.29 18.05 19.18 7.83 21.55 29.18
Table 3: We report SQuAD-$F_{1}$, ROUGE-L or SacreBLEU depending on the generation tasks. GPT-4 is evaluated in 0-shot, 3-shot and 6-shot setting.

### v3

@@ Abstract
In this paper, we compare general-purpose models, GPT-4-Turbo and Llama-3-8b, with special-purpose models—XLM-Roberta-large, mT5-large, and Llama-3-8b—that have been fine-tuned on specific tasks. We focus on seven classification and seven generation tasks to evaluate the performance of these models on Urdu language. Urdu has 70 million native speakers, yet it remains underrepresented in Natural Language Processing (NLP). Despite the frequent advancements in Large Language Models (LLMs), their performance in low-resource languages, including Urdu, still needs to be explored. We also conduct a human evaluation for the generation tasks and compare the results with the evaluations performed by GPT-4-Turbo, Llama-3-8b and Claude 3.5 Sonnet. We find that special-purpose models consistently outperform general-purpose models across various tasks. We also find that the evaluation done by GPT-4-Turbo for generation tasks aligns more closely with human evaluation compared to the evaluation the evaluation done by Llama-3-8b. This paper contributes to the NLP community by providing insights into the effectiveness of general and specific-purpose LLMs for low-resource languages.
@@
Samee Arif  and Abdul Hameed Azeemi  and Agha Ali Raza {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk Lahore University of Management Sciences
Awais Athar awais@ebi.ac.uk EMBL European Bioinformatics Institute
@@ 1 Introduction
    Generalists: General-purpose models capable of performing a wide variety of tasks. We will use GPT-4-Turbo (abbreviated as GPT) trained on dataset up to Dec 2023 and Llama-3-8b (abbreviated as Llama) as the generalist models.
@@
    Specialists: Special-purpose models fine-tuned to perform specific tasks. We use XLM-Roberta-large (abbreviated as XLM-R) (Conneau et al., 2020), mT5-large (abbreviated as mT5) (Xue et al., 2021), and a fine-tuned version of Llama-3-8b (abbreviated as Llama-FT) as the specialist models.
@@
Part-of-Speech Tagging Translation (ur-en)
Named-entity Recognition AI Assistant
@@
    We conduct a human evaluation of the outputs from the generation tasks and compare these results with the automated evaluations performed by GPT, Llama and Claude 3.5 Sonnet (abbreviated as Claude).
The code, model outputs, and the human and LLM evaluations are publicly available on GitHub [^1] https://github.com/sameearif/Generalists-vs-Specialists.
@@ 2 Related Work
In recent years there has been a growing interest in the performance of LLMs across various languages. The MEGA benchmark (Ahuja et al., 2023) evaluates 16 NLP datasets across 70 languages. They compare the performance of BLOOMZ, GPT models, and State of the Art (SOTA) non-autoregressive models. MEGAVERSE Ahuja et al. (2024) builds on top of the MEGA benchmark and evaluates the non-English capabilities of GPT-3.5-Turbo, GPT-4, PaLM2, Gemini-Pro, Mistral, Llama-2, and Gemma. IndicGenBench (Singh et al., 2024) evaluates LLMs on user-facing generation tasks across a set of 29 Indic languages. They perform evaluation on both proprietary and open-source LLMs including GP-3.5, GPT-4, PaLM-2, mT5, Gemma, BLOOM, and Llama. They cover the following tasks: cross-lingual summarization, machine translation, and cross-lingual question-answering. Zhao et al. (2024) conduct an evaluation of Llama’s response quality based on LLM-Eval (Zhang et al., 2023), a benchmark comprising instruction tasks from 17 categories. Mujadia et al. (2024) presents a comprehensive translation evaluation using Llama for English and Indian languages. They found that LLM-based evaluator achieves a comparable score with human judgement.
Khondaker et al. (2023a) presents a in-depth evaluaton of BLOOMZ, ChatGPT and specialist models AraT5 and MARBERTv2. They evaluated these models for Arabic on 44 distinct language understanding and generation tasks. They also present a comparison between the human evaluation and GPT-4 evaluation. Additionally, they observed that the English prompt works better than the Arabic prompt. Abdelali et al. (2024) provides a benchmark of LLMs against SOTA models for Arabic NLP. They evaluated GPT-3.5-Turbo, GPT-4, BLOOMZ, and Jais-13b-Chat across 33 tasks using 61 publicly available datasets, resulting in 330+ experimental setups. They observed that SOTA models generally outperform LLMs in zero-shot learning, larger models with few-shot learning techniques significantly reduce the performance gaps.
Existing multilingual NLP benchmarks often lack extensive language-specific evaluation and comparison against task-specific models. Tahir et al. (2024) address this gap by evaluating GPT-3.5-Turbo, Llama-2-7B-Chat, and Bloomz (3B and 7B) across 14 tasks using 15 Urdu datasets in a zero-shot setting, comparing their performance against task-specific models. Their findings reveal that task-specific models (Support Vector Machine (SVM), Decision Tree, and m-BERT, etc) generally outperform the mentioned LLMs in Urdu NLP tasks with zero-shot learning. However, they do not perform few-shot prompting, chain-of-thought prompting, or LLM fine-tuning.
@@ 3.1 Classification
Abuse Detection. We use the dataset by Akhter et al. (2020) which has 2,171 entries and the dataset by Amjad et al. (2022) which consists of 3,502 entries.
@@
Fake News Detection. We use the fake news dataset by Amjad et al. (2020) which is a mixture of real and translated data. It comprises 1,300 labeled news articles.
@@
Part-of-Speech (PoS) Tagging. For this task, we use the Universal Dependencies (Nivre et al., 2020) dataset, which consists of 5,130 sentences, annotated with a PoS tag for every word. For GPT and Llama, we wrap the word for which we want to predict the PoS tag in <hl> tags. The structure of the data is given in Figure 1.
@@
Named Entity Recognition (NER). We use the dataset [^4] https://huggingface.co/datasets/mirfan899/urdu-ner available on Hugging Face which consists of 33,748 sentences, annotated with a NER tag for every word. For GPT and Llama, we applied the same data structuring method as done in the PoS tagging task.
@@ 3.2 Generation
Given our resource constraints, we could not set the maximum sequence length for model training to the longest entries in each dataset. Instead, we filter the datasets to include only those entries that fell within a manageable maximum length of 2048. This approach allowed us to optimize the training process and ensure efficient use of computational resources while still maintaining a representative sample of the data.
Question-Answering. We use three datasets for question-answering: (1) UQA (Arif et al., 2024a), consisting of 88,829 answerable questions; (2) UQuAD [^5] https://github.com/ahsanfarooqui/UQuAD---Urdu-Question-Answer-Dataset/tree/main, containing 139 questions; and (3) Wiki-UQA [^6] https://huggingface.co/datasets/uqa/Wiki-UQA, a manually generated dataset from Wikipedia articles, comprising 210 questions.
@@
AI Assistant. We use UrduAssistant (Khalil, 2023) dataset which is translated to Urdu from English and has 67,017 prompts.
@@ 4.1 Experimental Design
We utilize a controlled test set consisting of min(1000, len(dataset["test"])) samples for each task except AI assistant, ensuring that the evaluation is both comprehensive and cost-efficient. For AI assistant we do human evaluation of 50 samples. We ensure that the test dataset is as balanced as possible, with an effort to achieve equal representation of different classes within each task. We use Macro-$F_{1}$ Score to evaluate all the classification tasks, SQuAD $F_{1}$ (Rajpurkar et al., 2018) Rajpurkar et al. (2016) to evaluate question-answering, SacreBLEU (Post, 2018) for paraphrasing, transliteration and translation, ROUGE-L (with word-level tokenization) Lin (2004) for summarization and Wins (the number of times a model is ranked one by the evaluator) for AI assistant evaluation.
@@
To evaluate the performance of the generalist models, we design a series of experiments. We use GPT and Llama as our generalist models. GPT is chosen due its top performance on the LMSYS chatbot arena (Chiang et al., 2024) as of March 1st, 2024 when we started our research. Our experimental setup is as follows: We conduct experiments under zero-shot, three-shot, and six-shot settings for both generation and classification tasks. The examples for the three-shot and six-shot scenarios are selected from the training dataset of each task. Specifically, the examples are carefully selected by a human expert to ensure a representative and balanced sample, avoiding any unintended bias in the selection process. We also conduct experiments with Chain-of-Thought (CoT) reasoning for classification tasks in the six-shot setting. We create CoT reasoning for the six selected examples from the training dataset. These examples along with their CoT is given as a few-shot prompt to the model. Figure 2 presents an example of generated CoT reasoning. For all the evaluations, the temperature and nucleus sampling for GPT-4 is set to 1.0, which is the default for the GPT API. For Llama, the temperature is set to 0.6 and nucleus sampling is set to 0.9, i.e., the values used in the original code-base for Llama [^8] https://github.com/meta-llama/llama3/blob/main/example_chat_completion.py.
@@ 4.2 Prompt Design
Khondaker et al. (2023b) observe in their ChatGPT evaluation for Arabic that an English prompt performs better than the Arabic one. Therefore, following this study, we decide to use English prompts for our evaluations of Urdu tasks as well. In the prompt templates, the following placeholders are used:
@@
4.3 Human Evaluation
We select a subset of size 50 (as done by Khondaker et al. (2023a) for Arabic) from the test dataset of each task. For human evaluation, two annotators (native Urdu speakers) are presented with anonymized outputs of Llama, Llama-FT, mT5 and GPT. They are asked to rank them from one to four based on the criteria in Appendix G, allowing multiple models to have the same rank. We prompt GPT, Llama and Claude with the same criteria, to assign a score to the outputs of the mentioned tasks, and then ranking is done based on this score. Since both GPT and Llama responses are being evaluated, we also include Claude as an additional evaluator because of the bias when LLMs assess their own outputs (Arif et al., 2024b). We compare ranking done by human annotators, and the three LLMs using Krippendorff’s alpha to determine the inter-rater reliability to determine the determine whether LLMs are effective evaluators for Urdu tasks.
@@ 5.1 Classification
GPT Llama XLM-R Llama-FT
Task 0 3 6 CoT 0 3 6 CoT
Sentiment Analysis 90.98 91.17 94.90 94.60 87.88 60.62 59.71 57.54 92.90 95.30
Abuse Detection 86.27 89.01 88.71 87.62 44.73 65.85 71.64 75.82 90.92 89.15
Sarcasm Detection 58.17 69.18 66.56 65.47 44.94 34.25 50.67 49.75 84.37 81.48
Fake News Detection 78.88 75.95 78.14 76.45 66.36 63.48 71.45 40.98 84.99 72.14
Topic Classification 76.05 73.54 74.57 70.43 54.08 53.02 53.68 49.24 84.49 84.74
PoS Tagging 53.31 51.51 54.17 54.61 25.80 34.32 33.85 41.62 65.41 67.55
NER Tagging 61.96 62.18 62.98 63.95 42.13 40.80 45.29 53.96 70.41 90.90
Table 2: We report Macro-$F_{1}$ score for each classification task. GPT-4-Turbo and Llama-3-8b is evaluated in 0-shot, 3-shot, 6-shot and 6-shot with Chain-of-Thought settings. FT stands for fine-tuned.
We present the evaluation of the generalists (GPT, Llama) and the specialists (Llama-FT, XLM-R) for the classification tasks in Table 2. We observe that Llama-FT (fine-tuned) model achieves the highest scores in four tasks (sentiment analysis, topic classification, PoS tagging, and NER tagging), while XLM-R outperforms other models in three tasks (abuse detection, sarcasm detection, and fake news detection). GPT does not perform better than Llama-FT and XLM-R. For certain tasks like NER and PoS tagging, CoT reasoning leads to a better Macro-$F_{1}$ score. We now discuss the performance on each classification task in detail.
Sentiment Analysis. For sentiment analysis, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 95.30, outperforming XLM-R and GPT. GPT’s performance improves with more shots, reaching a Macro-$F_{1}$ of 94.90 with 6-shot learning. The CoT setting provides a slight increase to 94.60. However the Macro-$F_{1}$ for Llama decreases with few-shot and CoT prompting. The small difference in the Macro-$F_{1}$ between Llama-FT and GPT indicates the effectiveness of GPT for Urdu sentiment analysis without requiring task-specific fine-tuning.
@@
NER Tagging. For NER tagging, Llama-FT achieves the highest score with a Macro-$F_{1}$ of 90.90, significantly outperforming XLM-R and GPT. The highest score achieved by GPT is 63.95, and by Llama is 53.96, both using CoT reasoning. We uncover the differences in accuracy for various entities for GPT and Llama-FT in Figure 8. We observe a drop in the recognition accuracy of Person entities by GPT, with 95 correctly recognized compared to 139 by Llama-FT. We notice a similar trend for Organization entities, with 70 classified correctly compared to 134 by Llama-FT. This suggests that NER fine-tuning on a specialized Urdu dataset improves the model’s ability to recognize Person and Organization entities in Urdu text.
Abuse Detection. In the task of abuse detection, XLM-R leads with a Macro-$F_{1}$ of 90.92, followed closely by Llama-FT at 89.15. GPT achieves its highest performance of 89.01 with 3-shot prompting, while Llama reaches a peak of 75.82 using CoT reasoning.
Sarcasm Detection. XLM-R performs best in sarcasm detection with a Macro-$F_{1}$ score of 84.37. Llama-FT also shows strong performance with 81.48. GPT achieves its best performance at 69.18 with 3-shot prompting, while Llama performs best at 50.67 with 6-shot prompting. The large difference between the Macro-$F_{1}$ scores of generalist and specialist models indicates that sarcasm detection in Urdu can be challenging without task-specific fine-tuning.
Fake News Detection. For fake news detection, XLM-R achieves the highest score with a Macro-$F_{1}$ of 84.99, while Llama-FT follows with 72.14. GPT’s performance is relatively consistent across different shot settings, with its highest score being 78.88 in the 0-shot setting. Llama’s highest score is 71.45 at 6-shot setting. The large margin by XLM-R indicates its effectiveness in discerning fake news.
@@
GPT Llama mT5 Llama-FT
Task Metric 0 3 6 0 3 6
Question-Answering SQuAD-$F_{1}$ 66.28 72.55 72.78 69.89 71.65 71.62 69.66 70.42
Summarization ROUGE-L 22.54 23.35 23.34 25.76 25.84 26.17 30.72 31.01
Paraphrasing SacreBLEU 3.59 3.92 4.01 2.42 6.40 6.11 11.98 10.17
Transliteration SacreBLEU 30.93 32.38 32.09 12.37 18.51 21.35 40.23 37.95
Translation (en-ur) SacreBLEU 12.07 12.62 11.59 5.14 6.56 5.83 18.35 14.63
Translation (ur-en) SacreBLEU 16.29 18.05 19.18 7.83 13.11 12.32 21.55 29.18
Table 3: We report SQuAD-$F_{1}$, ROUGE-L or SacreBLEU depending on the generation tasks. GPT-4-Turbo and Llama-3-8b is evaluated in 0-shot, 3-shot and 6-shot setting.
@@
PoS Tagging. In PoS tagging, Llama-FT leads with a Macro-$F_{1}$ of 67.55, followed by XLM-R at 65.41. GPT’s highest score is 53.31 in the 0-shot setting and Llama’s highest score is 41.62 with CoT reasoning. To understand the significant difference in Macro-$F_{1}$ between GPT and Llama-FT, we study the individual class performance (Figure 9). We find that GPT struggles with correctly tagging proper nouns, pronouns, and auxiliaries, while Llama-FT is able to identify most of them correctly, suggesting that task-specific fine-tuning improves tagging performance for these parts of speech.